# NeuroRVQ: Multi-Scale EEG Tokenization for Generative Large Brainwave Models

### v1

@@ 5.3 Downstream Performance
We evaluated the downstream task performance of fine-tuned NeuroRVQ in comparison to other fine-tuned LBMs, namely NeuroGPT (Cui et al., 2024), CBraMod (Wang et al., 2025), LaBraM (Jiang et al., 2024), EEGPT and BIOT (Yang et al., 2023). [^1] BIOT could not be tested on Sleep since the benchmark electrodes are missing from the pre-trained model. All models were evaluated in four downstream classification tasks (Lee et al., 2025): (a) Motor, a four-class movement classification on High Gamma (Schirrmeister et al., 2017) dataset; (b) Memory, a binary working memory classification on Pavlov et al. (2022); (c) Sleep, a 6-class sleep stage classification on Physionet’s Sleep-EDF (Kemp et al., 2000); and (d) Eyes, eyes open vs. closed classification on Physionet’s Motor (Schalk et al., 2004).These tasks capture a diverse range of BCI paradigms and the datasets were specifically chosen for their minimal spurious artifacts, reducing the likelihood of spurious performance during training (Lee et al., 2025).
@@
Table 4: Comparison of classification balanced accuracy between NeuroRVQ and other LBMs, finetuned on the same datasets using identical training settings ( denotes early stopping). Bold and underlined values indicate best performance and next-best performance respectively (per task or overall). Size denotes the number of parameters for the backbone foundation models excluding the classification heads (see Appendix G).
Model Motor Memory Sleep Eyes Mean Size
@@
Table 4 shows that, when full-finetuned, NeuroRVQ achieves the best or next-best performance in each task, and the best overall performance, measured as the mean accuracy across all tasks. NeuroRVQ has 5.9M parameters in the backbone model and approximately 3,000 parameters in the classification heads. Yet, it achieves higher mean accuracy than NeuroGPT, which has 79.5M parameters in its backbone; and it also outperforms CBraMod, which has similar backbone parameters to NeuroRVQ but much larger classification heads (approximately 50M parameters; see Appendix G for a detailed comparison of classification head sizes). This combination of strong generalization and compact size highlights the effectiveness of our NeuroRVQ tokenizer.
@@ 7 Conclusion
In this work, we introduce NeuroRVQ, an LBM that leverages a Residual Vector Quantization (RVQ) codebook-based tokenizer. Our main contribution is the design and proper training of the NeuroRVQ tokenizer, which captures low- and high-frequency components of EEG signals, achieves state-of-the-art reconstruction performance and facilitates the development of more effective LBMs. We further showed that NeuroRVQ outperforms existing pre-trained LBMs on various downstream tasks. In summary, NeuroRVQ advances the development of efficient LBMs by providing a high-fidelity tokenizer. The principles underlying NeuroRVQ’s tokenizer can be extended to other biosignals, paving the way for broader applications of foundation models in the biosignal space.

### v2

@@ 5.3 Downstream Performance
We evaluated the downstream task performance of fine-tuned NeuroRVQ in comparison to other fine-tuned LBMs, namely NeuroGPT (Cui et al., 2024), CBraMod (Wang et al., 2025), LaBraM (Jiang et al., 2024), EEGPT (Wang et al., 2024), BIOT (Yang et al., 2023) [^1] BIOT could not be tested on Sleep since the benchmark electrodes are missing from the pre-trained model. and MIRepNet (Liu et al., 2025) [^2] As a motor-related foundation model, MIRepNet was tested only on the motor task. All models were evaluated in four downstream classification tasks (Lee et al., 2025): (a) Motor, a four-class movement classification on High Gamma (Schirrmeister et al., 2017) dataset; (b) Memory, a binary working memory classification on Pavlov et al. (2022); (c) Sleep, a 6-class sleep stage classification on Physionet’s Sleep-EDF (Kemp et al., 2000); and (d) Eyes, eyes open vs. closed classification on Physionet’s Motor (Schalk et al., 2004).These tasks capture a diverse range of BCI paradigms and the datasets were specifically chosen for their minimal spurious artifacts, reducing the likelihood of spurious performance during training (Lee et al., 2025).
@@
Table 4: Comparison of classification balanced accuracy between NeuroRVQ and other LBMs, finetuned on the same datasets using identical training settings ( denotes early stopping). Bold and underlined values indicate best performance and next-best performance respectively (per task or overall).
Model Motor ERP Memory Sleep Eyes Mean
@@
[…]
Table 4 shows that, when full-finetuned, NeuroRVQ achieves the best or next-best performance in each task, and the best overall performance, measured as the mean accuracy across all tasks. NeuroRVQ has 5.9M parameters in the backbone model and approximately 3,000 parameters in the classification heads (see Appendix G). Yet, it achieves higher mean accuracy than NeuroGPT, which has 79.5M parameters in its backbone; and it also outperforms CBraMod, which has similar backbone parameters to NeuroRVQ but much larger classification heads (approximately 50M parameters; see Appendix G for a detailed comparison of classification head sizes). This combination of strong generalization and compact size highlights the effectiveness of our NeuroRVQ tokenizer.
@@ 7 Conclusion
In this work, we introduce NeuroRVQ, an LBM that leverages a Residual Vector Quantization (RVQ) codebook-based tokenizer. Our main contribution is the design and proper training of the NeuroRVQ tokenizer, which captures low- and high-frequency components of EEG signals, achieves state-of-the-art reconstruction performance and facilitates the development of more effective LBMs. We further showed that NeuroRVQ outperforms existing pre-trained LBMs on various downstream tasks. In summary, NeuroRVQ advances the development of efficient LBMs by providing a high-fidelity tokenizer. The principles underlying NeuroRVQ’s tokenizer can be extended to other biosignals (see Appendix H), paving the way for broader applications of foundation models in the biosignal space.
@@ Appendix G Classification Heads: Architecture & Size
Appendix H Tokenization of Biosignal Modalities
The principles underlying NeuroRVQ’s tokenizer apply to other biosignals, enabling broader applications of foundation models in the biosignal domain. In this section, we present signal reconstruction results on two additional modalities: electromyography (EMG) and electrocardiogram (ECG).
For ECG, we used the PTB-XL dataset (Wagner et al., 2022), resampled at 200Hz. The NeuroRVQ tokenizer was trained using the settings detailed in Appendix B. Training was conducted for 100 epochs with $S=4$ temporal branches and 4 $\mathcal{R}$ RVQ codebooks, each consisting of 8 single codebooks $\mathcal{V}{i}\in\mathbb{R}^{8192\times 128}$, with $\lambda_{circle}=0.4$. Figure 7 shows signal reconstruction for an ECG signal, demonstrating excellent fidelity.
For EMG, we used the emg2pose dataset (Salter et al., 2024), resampled at 1000Hz. The NeuroRVQ tokenizer used the same configuration as above, except that each of the 4 $\mathcal{R}$ RVQ codebooks contained 16 single codebooks $\mathcal{V}_{i}\in\mathbb{R}^{8192\times 128}$ and the kernel sizes approprietely adapted as detailed in Table 10. Figure 8 shows signal reconstruction for an EMG signal, showing high accuracy in preserving its original features.
Table 10: Configuration of the multi-scale temporal encoder used in the EMG tokenizer. Here, $x\rightarrow y$ indicates values for the first ($x$) and second ($y$) sequence.
Branch № Filters Kernel Padding Pooling
1 8 $\rightarrow$ 8 (1, 51) $\rightarrow$ (1, 25) (0, 25) $\rightarrow$ (0, 12) (1, 2) $\rightarrow$ (1, 4)
2 8 $\rightarrow$ 8 (1, 17) $\rightarrow$ (1, 9) (0,   8) $\rightarrow$ (0, 4) (1, 2) $\rightarrow$ (1, 4)
3 8 $\rightarrow$ 8 (1,   8) $\rightarrow$ (1, 4) (0,   4) $\rightarrow$ (0, 2) (1, 2) $\rightarrow$ (1, 4)
Figure 7: Per-Band Analysis of Reconstructed ECG signal from NeuroRVQ codebook-based tokenizer. Blue lines denote the input EMG signal and orange the reconstructed ECG signal for NeuroRVQ.
Figure 8: Per-Band Analysis of Reconstructed EMG signal from NeuroRVQ codebook-based tokenizer. Blue lines denote the input EMG signal and orange the reconstructed EMG signal for NeuroRVQ.